OpenStudy (kittiwitti1):

SOLVED - READ BOTTOM QUESTION. *** What is the work done from B to A if n=4 mol? A: -18000 J I got -12000 J... HUGE difference.

OpenStudy (kittiwitti1):

Information (V is volume and P is pressure - originally given in kPa)\[V_B=10 m^3,~P_B=1\times10^3 Pa\]\[V_A=4m^3,~P_A=5\times10^3Pa\]*** Formulas\[\Delta U=\frac{3}{2}P\Delta V=\frac{3}{2}nRT\]\[P\Delta V=nRT\]

OpenStudy (agent0smith):

Can't you solve the very last equation for T? Then plug it into the prior equation.

OpenStudy (kittiwitti1):

For T?

OpenStudy (kittiwitti1):

But doesn't\[W=P\Delta V\]

OpenStudy (kittiwitti1):

o.o?

OpenStudy (agent0smith):

^wouldn't that be for constant pressure...? Idk, i'm too tired to think about it

OpenStudy (anonymous):

"PΔV=nRT" that equation is definitely wrong :-/

OpenStudy (anonymous):

you have only provided the intial and final states..and not mentioned the type of process work done depends upon the PATH taken (the process u used to go from state A to state B).. so unless u mention that, u can't calculate work done!

OpenStudy (anonymous):

\[W = \int\limits_{A}^{B}PdV\]

OpenStudy (kittiwitti1):

@Mashy this is what we were taught...

OpenStudy (kittiwitti1):

I believe the process is .... adiabatic?

OpenStudy (anonymous):

u believe? :P.. or you know? cause different processes can give u different answers.. do you know what is the formula for the work done in case of adiabatic process ?

OpenStudy (kittiwitti1):

Well it's a triangle, and B>A is the hypotenuse. Since it's a pressure-volume graph work would be PV or area under the curve.

OpenStudy (anonymous):

yea.. work done is area under the curve.. the graph is given to you? mentioned in the question itself?

OpenStudy (kittiwitti1):

It is given but I cannot post it. B is the bottom corner....

OpenStudy (kittiwitti1):

That's the best I can do....

OpenStudy (anonymous):

u did not mention this :P.. without mentioning the graph how do u expect anyone to solve it? :D :D The graph itself indicates what process it is.. (its not adiabatic.. not important now what the process is anyways.. cause the graph is known) ok .. lemme see if i can draw it better!

OpenStudy (raffle_snaffle):

This looks exactly like one of my hw problems.....

OpenStudy (raffle_snaffle):

the work done from B to A is w=1/2*P*V

OpenStudy (raffle_snaffle):

No work is being done from A to C

OpenStudy (anonymous):

 now the area equals work done under curve = shaded region !